# A faster FPRAS for #NFA

### 20 Dec 2023 v1

@@ 2     $\eta^{\prime}\leftarrow\eta/4n$ $\beta^{\prime}\leftarrow(1+\beta)^{\ell-1}-1$ if $\ell=0$ then
1 $\beta\leftarrow\frac{\varepsilon}{4n^{2}}$ $\eta\leftarrow\frac{\delta}{2\cdot n\cdot m}$ $\mathtt{ns}\leftarrow\frac{4096\cdot e\cdot n^{4}}{\epsilon^{2}}\log\left(\frac{4096\cdot m^{2}\cdot n^{2}\cdot\log(\varepsilon^{-2})}{\delta}\right)$ $\mathtt{xns}\leftarrow\mathtt{ns}\cdot 12\cdot(1-\frac{2}{3e^{2}})^{-1}\cdot\log(8/\eta)$ Construct the labeled directed acyclic graph $\mathcal{A}_{\textsf{unroll}}$ from $\mathcal{A}$ input $n$ for $\ell\in\{0,\ldots,n\}\text{ and }q\in Q$ do
@@
3          $\big{(}\mathsf{N}({q}^{\ell}),\mathsf{S}({q}^{\ell})\big{)}\leftarrow(q\in F)$ $\big{(}1,\{\lambda\}\big{)}$ : $\big{(}0,\varnothing)$
@@
5          $\{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1},\ldots,p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}\}\leftarrow\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})$ $\beta^{\prime}=(1+\beta)^{\ell-1}-1$ $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leftarrow\texttt{AppUnion}_{\beta,\eta/2}\bigg{(}\beta^{\prime},$ $(\mathsf{S}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1}}^{\ell-1}),\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1}}^{\ell-1})),\ldots,(\mathsf{S}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}}^{\ell-1}),\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}}^{\ell-1})),$ $\mathcal{A}_{\textsf{unroll}}\bigg{)}$   // The last parameter $\mathcal{A}_{\textsf{unroll}}$ is a proxy for the membership oracles required by AppUnion
6
7    $\mathsf{N}({q}^{\ell})\leftarrow\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ $\mathsf{S}({q}^{\ell})\leftarrow\varnothing$ repeat $\mathtt{xns}$ times:
@@ 10                    $\mathsf{S}({q}^{\ell})\leftarrow\mathsf{S}({q}^{\ell})\boldsymbol{\!\cdot}\!\mathsf{append}(w)$
Algorithm 4 Algorithm for Estimating $|\mathcal{L}(\mathcal{A})_{n}|$
The main algorithm. Algorithm 4 first constructs (in Line 4) the labeled directed acyclic graph $\mathcal{A}_{\textsf{unroll}}$. The algorithm then goes over all the states and all levels $0\leq\ell\leq n$ to construct the sampled sets and the size estimates inductively. In Lines 4-4, the algorithm caters for the base case ($\ell=0$). In the inductive case (Lines 4-4) the algorithm computes the estimate $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ and $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$, where $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the estimate of the size of $\cup_{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},i}\in\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},i}}^{\ell-1})$. The size estimates of, and a sample subsets of the sets $\{\mathcal{L}({q}^{\ell-1})\}_{q\in Q}$ are available inductively at this point. Further, membership in these languages can be easily checked. As a result, the algorithm uses the subroutine AppUnion to compute the size of the union (in Line 4). Once the size estimates $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$ and $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ are determined, the algorithm constructs the sampled subset of $\mathcal{L}({q}^{\ell})$ by making $\mathsf{xns}$ calls to the subroutine sample (Lines 4-4). Theorem 3 presents the desired correctness statement of our FPRAS  Algorithm 4, and its proof is presented in Section 6.
Theorem 3.
@@
${\mathsf{Pr}}\!\left[\frac{|L(F^{n})|}{1+\epsilon}\leq\mathsf{Est}\leq(1+\epsilon)|L(F^{n})|\right]\geq 1-\delta$
@@
Moreover, the algorithm has time complexity $\widetilde{O}(m^{4}n^{11}\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$, where the tilde hides for polynomial factors of $\log(m+n)$.
4. Correctness of Algorithm 2
@@
The proof is in the same lines as that of Karp and Luby (Karp and Luby, 1985); for completeness, we present a full proof.
Proof.
@@
Let $\mathsf{Fail}$ be the event that the output of the algorithm is not within $\frac{|\cup_{i=1}^{k}T_{i}|}{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}$ and $|\cup_{i=1}^{k}T_{i}|{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}$. Clearly, we need to upper bound $\Pr[\mathsf{Fail}]$ by $\delta$.
@@
Note that the algorithm in each of the $t$ runs of the loop (lines 2-2) tries to draw an element from some $S_{i}$. So if we assume that the size of the sets $S_{i}$ for all $i$ is more than $t$, then the condition is the if (in line 2 will be satisfied and in that case the else clause in line 2 to 2 is redundant.
@@
We will prove the correctness of  Algorithm 2 in two parts. In the first part we will prove that if the size of the sets $S_{i}$ is greater than $t$ then $\Pr[\mathsf{Fail}]\leq\frac{\delta}{2}$. In the second part we will prove that if the sets $S_{i}$ has size thresh (much smaller than $t$) then the probability that the algorithm will ever need the else clause in line 2 to 2 is less that $\delta/2$. The two parts combined proves that if the sets $S_{i}$s are of size threshthen $\Pr[\mathsf{Fail}]\leq\delta$ which is what is claimed in the theorem.
@@
For all $\sigma\in\cup_{i=1}^{k}T_{i}$ let $L(\sigma)$ denote the smallest index $i$ such that $\sigma\in T_{i}$. In other words, $\sigma$ is in $T_{L(\sigma)}$ and for all $i<L(\sigma)$, $\sigma\not\in T_{i}$. Observe $\mathcal{U}_{1}:=\left\{(\sigma,L(\sigma)\ \mid\ \sigma\in\cup_{i=1}^{k}T_{i}\right\}$ has size exactly $\left|\cup_{i=1}^{k}T_{i}\right|$. Furthermore, $\mathcal{U}_{2}:=\left\{(\sigma,i)\ \mid\ \sigma\in T_{i}\right\}$ has size $\sum_{i=1}^{k}|T_{i}|$.
Part 1: If $\forall i$, $|S_{i}|\geq t$ then $\Pr[\mathsf{Fail}]\leq\frac{\delta}{2}$.
@@
Note that in this case the condition in the if clause in line 2 is always satisfied. For a run of the loop (lines 2-2), consider the lines 2 to 2. We say the pair $(\sigma,i)$ is sampled if in one of the iteration of the loop $i$ is sampled in line 2 and then in that same iteration $\sigma$ is obtained from $S_{i}.\mathsf{deque()}$ in the line 2.
For any $i_{0}\in[t]$ and $\sigma_{0}\in S_{i_{0}}$, we have

### 08 Apr 2024 v2

@@ 3.3. Main Algorithm
1 $\beta\leftarrow\frac{\varepsilon}{4n^{2}};\eta\leftarrow\frac{\delta}{2\cdot n\cdot m}$ $\mathtt{ns}\leftarrow\frac{4096\cdot e\cdot n^{4}}{\epsilon^{2}}\log\left(\frac{4096\cdot m^{2}\cdot n^{2}\cdot\log(\varepsilon^{-2})}{\delta}\right)$ $\mathtt{xns}\leftarrow\mathtt{ns}\cdot 12\cdot(1-\frac{2}{3e^{2}})^{-1}\cdot\log(8/\eta)$ Construct the labeled directed acyclic graph $\mathcal{A}_{\textsf{unroll}}$ from $\mathcal{A}$ for $\ell\in\{0,\ldots,n\}\text{ and }q\in Q$ do
@@
3          if $q=I$ then
4               $\big{(}\mathsf{N}({q}^{\ell}),\mathsf{S}({q}^{\ell})\big{)}\leftarrow(1,{\lambda})$
5         else
6               $\big{(}\mathsf{N}({q}^{\ell}),\mathsf{S}({q}^{\ell})\big{)}\leftarrow\big{(}0,\varnothing)$
7    else
@@ 6               $\big{(}\mathsf{N}({q}^{\ell}),\mathsf{S}({q}^{\ell})\big{)}\leftarrow\big{(}0,\varnothing)$
9               let $\{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1},\ldots,p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}\}$ be the set $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})$ $\beta^{\prime}\leftarrow(1+\beta)^{\ell-1}-1$ $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leftarrow\texttt{AppUnion}_{\beta,\frac{\eta}{2}\cdot(1-\frac{1}{2^{n+1}})}\bigg{(}\beta^{\prime},(\mathsf{S}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1}}^{\ell-1}),\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},1}}^{\ell-1})),\ldots,(\mathsf{S}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}}^{\ell-1}),\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},k_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}}^{\ell-1})),\mathcal{A}_{\textsf{unroll}}\bigg{)}$
10         w.p.$1-\frac{\eta}{2^{n}}$ :
11               $\mathsf{N}({q}^{\ell})\leftarrow\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}+\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$
12         w.p.$\frac{\eta}{2^{n}}$ :
[…]
14          $\mathsf{S}({q}^{\ell})\leftarrow\varnothing$ repeat $\mathtt{xns}$ times:
@@
Algorithm 3 Algorithm for estimating $|\mathcal{L}(\mathcal{A}_{n})|$
Algorithm 3 first constructs (in Line 3) the labeled directed acyclic graph $\mathcal{A}_{\textsf{unroll}}$. The algorithm then goes over all the states and all levels $0\leq\ell\leq n$ to construct the sampled sets and the size estimates inductively. In Lines 3-3, the algorithm caters for the base case ($\ell=0$). In the inductive case (Lines 3-3) the algorithm computes the estimate $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ and $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$, where $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the estimate of the size of $\cup_{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},i}\in\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},i}}^{\ell-1})$. The size estimates of, and a sample subsets of the sets $\{\mathcal{L}({q}^{\ell-1})\}_{q\in Q}$ are available inductively at this point. Further, membership in these languages can be easily checked. As a result, the algorithm uses the subroutine AppUnion to compute the size of the union (in Line 3), with $\mathcal{A}_{\textsf{unroll}}$ as the membership oracle as before. Once the size estimates $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}$ and $\textsf{sz}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}$ are determined, the algorithm constructs the sampled subset of $\mathcal{L}({q}^{\ell})$ by making $\mathsf{xns}$ calls to the subroutine sample (Lines 3-3); if fewer than $\mathtt{ns}$ many strings were sampled, the algorithm simply pads one fixed string from $\mathcal{L}({q}^{\ell})$ so that $\mathsf{S}({q}^{\ell})$ eventually has size $\mathtt{ns}$ (see Lines 3-3). Theorem 3 presents the desired correctness statement of our FPRAS  Algorithm 3, and its proof is presented in Section 4.3.
Theorem 3 ().
@@
${\mathsf{Pr}}\!\left[\frac{|\mathcal{L}(\mathcal{A}_{n})|}{1+\epsilon}\leq\mathsf{Est}\leq(1+\epsilon)|\mathcal{L}(\mathcal{A}_{n})|\right]\geq 1-\delta$
@@
Moreover, the algorithm has time complexity $\widetilde{O}((m^{2}n^{10}+m^{3}n^{6})\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$, where the tilde hides for polynomial factors of $\log(m+n)$.
4. Technical Analysis
We now present the detailed technical analyses of the algorithms presented in Section 3.
4.1. Correctness of Algorithm 1
@@
The proof is in the same lines as that of Karp and Luby (Karp and Luby, 1985). For the sake of completeness, we present a full proof.
Proof (of Theorem 1).
@@
Let $\mathsf{Fail}$ be the event that the output of the algorithm is not within $\frac{|\cup_{i=1}^{k}T_{i}|}{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}$ and $|\cup_{i=1}^{k}T_{i}|{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}$. We will show that ${\mathsf{Pr}}\!\left[\mathsf{Fail};{{\mathsf{U}}_{1,\ldots,k}}\right]$ is upper-bounded by $\delta$.
@@
Note that the algorithm in each of the $t$ runs of the loop (Lines 1-1) tries to draw an element from some $S_{i}$. So if we assume that the size of the sets $S_{i}$ for all $i$ is more than $t$, then the condition in the if (in Line 1) will be satisfied and in that case the else clause in Line 1 is redundant.
@@
We will prove the correctness of  Algorithm 1 in two parts. In the first part we will prove that if the size of the sets $S_{i}$ is greater than $t$ then ${\mathsf{Pr}}\!\left[\mathsf{Fail};{{\mathsf{U}}_{1,\ldots,k}}\right]\leq\frac{\delta}{2}$. In the second part we will prove that if the sets $S_{i}$ has size thresh (much smaller than $t$) then the probability that the algorithm will ever need the else clause in Line 1 is less that $\delta/2$. The two parts combined proves that if the sets $S_{i}$s are of size thresh, then ${\mathsf{Pr}}\!\left[\mathsf{Fail};{{\mathsf{U}}_{1,\ldots,k}}\right]\leq\delta$ which is what is claimed in the theorem.
@@
For all $\sigma\in\cup_{i=1}^{k}T_{i}$ let $L(\sigma)$ denote the smallest index $i$ such that $\sigma\in T_{i}$. In other words, $\sigma$ is in $T_{L(\sigma)}$ and for all $i<L(\sigma)$, $\sigma\not\in T_{i}$. Observe $\mathcal{U}_{\textsf{unique}}:=\left\{(\sigma,L(\sigma))\ \mid\ \sigma\in\cup_{i=1}^{k}T_{i}\right\}$ has size exactly $\left|\cup_{i=1}^{k}T_{i}\right|$. Furthermore, $\mathcal{U}_{\textsf{multiple}}:=\left\{(\sigma,i)\ \mid\ \sigma\in T_{i}\right\}$ has size $\sum_{i=1}^{k}|T_{i}|$.
@@
Part 1: If $\forall i$, $|S_{i}|\geq t$ then ${\mathsf{Pr}}\!\left[\mathsf{Fail};{{\mathsf{U}}_{1,\ldots,k}}\right]\leq\frac{\delta}{2}$. If $|S_{i}|\leq t$ for all $i$, then the condition in the if clause in Line 1 is always satisfied. For a run of the loop (Lines 1-1), consider the Lines 1-1. We say the pair $(\sigma,i)$ is sampled in round $j$ if in the $j^{\text{th}}$ iteration of the loop $i$ is sampled in Line 1 and then in that same iteration $\sigma$ is obtained from $S_{i}.\mathsf{dequeue()}$ in the Line 1.
For any $i_{0}\in\{1,\ldots,k\},\sigma_{0}\in S_{i_{0}}$ and $j_{0}\in\{1,\ldots,t\}$, we have
[…]
Since the choice of the set to pick $i_{0}$ is independent of the distributions of any of the $S_{i}$’s, we have,
[…]
Next observe that the second term in Equation (1) is ${\mathsf{Pr}}\!\left[\sigma_{0}\mbox{ is sampled in round $j_{0}$ from }S_{i_{0}};{{\mathsf{U}}_{1,\ldots,k}}\right]$. This term, in turn, can be computed by considering the disjoint union of the events corresponding to how many times $i_{0}$ has been chosen so far: ${\mathsf{Pr}}\!\left[\sigma_{0}\mbox{ is sampled in round $j_{0}$ from }S_{i_{0}};{{\mathsf{U}}_{1,\ldots,k}}\right]$
$=$ $\sum\limits_{m=1}^{t}{\mathsf{Pr}}\!\left[\begin{aligned} i_{0}\text{ was sampled $m-1$ times before round }j_{0}\\ \text{ and }\sigma_{0}\mbox{ is sampled in round $j_{0}$ from }S_{i_{0}}\end{aligned};{\mathsf{U}}_{1,\ldots,k}\right]=\sum\limits_{m=1}^{t}{\mathsf{Pr}}\!\left[S_{i_{0}}[m]=\sigma_{0};{{\mathsf{U}}_{1,\ldots,k}}\right]$
$=$ $\sum\limits_{\omega}\sum\limits_{m=1}^{t}{\mathsf{Pr}}\!\left[\omega[m]=\sigma_{0}\mid S_{i_{0}}=\omega\right]\cdot{\mathsf{Pr}}\!\left[U_{i_{0}}=\omega\right]=\sum\limits_{\omega}{\mathsf{Pr}}\!\left[U_{i_{0}}=\omega\text{ and }\omega[m]=\sigma_{0}\right]=\frac{1}{|T_{i_{0}}|}$
Therefore, $\frac{1}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq{\mathsf{Pr}}\!\left[(\sigma_{0},i_{0})\mbox{ is sampled in round }j_{0};{{\mathsf{U}}_{1,\ldots,k}}\right]\leq(1+\epsilon_{\textsf{sz}})\frac{1}{\sum_{i=1}^{k}\textsf{sz}_{i}}.$
For any $i_{0}\in[t]$ and $\sigma_{0}\in S_{i_{0}}$, we have ${\mathsf{Pr}}\!\left[(\sigma_{0},i_{0})\mbox{ is sampled };{{\mathsf{U}}_{1,\ldots,k}}\right]$
[…]
$=$ $\frac{\textsf{sz}_{i_{0}}}{\sum_{i}^{k}\textsf{sz}_{i}}\times\frac{1}{|T_{i_{0}}|}$
Probability that $i_{0}$ is sampled in Line 1 is independent of the distributions of the samples. The last equality is because we have assumed $\forall i$, $|S_{i}|\geq t$, so the if condition in Line 1 is always satisfied. Hence once $i_{0}$ is sampled in Line 1 probability that $\sigma_{0}$ is obtained is exactly $1/|T_{i_{0}}|$, since the set $S_{i_{0}}$ contains elements sampled uniformly (with replacement) from $T_{i_{0}}$. Therefore, $\frac{1}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq{\mathsf{Pr}}\!\left[\sigma_{0},i_{0})\mbox{ is sampled };{{\mathsf{U}}_{1,\ldots,k}}\right]\leq(1+\epsilon_{\textsf{sz}})\frac{1}{\sum_{i=1}^{k}\textsf{sz}_{i}}.$
Let $Y_{j}$ be the random variable denoting whether the counter $Y$ is increased in the $j$th iteration of the loop. Note that $Y_{j}=1$ if the pair $(\sigma_{0},i_{0})$ sampled in the $j$th iteration is in $\mathcal{U}_{\textsf{unique}}$. This is because in Line 1 it is checked if the sampled pair is in $\mathcal{U}_{\textsf{unique}}$.
So, $\mathbb{E}[Y_{j};{\mathsf{U}}_{1,\ldots,k}]=\sum_{(\sigma,i)\in\mathcal{U}_{\textsf{unique}}}{\mathsf{Pr}}\!\left[(\sigma,i)\mbox{ is sampled in round }j;{{\mathsf{U}}_{1,\ldots,k}}\right]$; here the ‘;’ in the expectation has an analogous definition to the case when used for probability. Thus,
$\frac{|\mathcal{U}_{\textsf{unique}}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq\mathbb{E}[Y_{j};{\mathsf{U}}_{1,\ldots,k}]\leq(1+\epsilon_{\textsf{sz}})\frac{|\mathcal{U}_{\textsf{unique}}|}{\sum_{i=1}^{k}\textsf{sz}_{i}}$
Note that $Y=\sum_{i}^{t}Y_{i}$. Thus at the end of the algorithm the expected value of $Y$ is between $\frac{t|\mathcal{U}_{\textsf{unique}}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}$ and $(1+\epsilon_{\textsf{sz}})^{2}\frac{t|\mathcal{U}_{\textsf{unique}}|}{\sum_{i=1}^{k}\textsf{sz}_{i}}$. Now by Chernoff bound we have,
${\mathsf{Pr}}\!\left[Y\not\in(1\pm\frac{\epsilon}{2})\mathbb{E}[Y_{j};{\mathsf{U}}_{1,\ldots,k}];{{\mathsf{U}}_{1,\ldots,k}}\right]\leq 2exp(-\frac{\epsilon^{2}\mathbb{E}[Y_{j};{\mathsf{U}}_{1,\ldots,k}]}{12}).$
We also know that,
$\mathbb{E}[Y;{\mathsf{U}}_{1,\ldots,k}]\geq$ $\frac{t|\mathcal{U}_{\textsf{unique}}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\geq\frac{t\max_{i}|T_{i}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\geq\frac{t\max_{i}\textsf{sz}_{i}}{(1+\epsilon_{\textsf{sz}})^{2}\sum_{i=1}^{k}\textsf{sz}_{i}}\geq 12\log\left(\frac{4}{\delta}\right)$
The last inequality is due to the setting of the parameter $t$. Thus, by Chernoff bound, probability that the output of the algorithm is not between $(1-\frac{\epsilon}{2})\frac{|\mathcal{U}_{\textsf{unique}}|}{(1+\epsilon_{\textsf{sz}})}$ and $(1+\frac{\epsilon}{2})(1+\epsilon_{\textsf{sz}})|\mathcal{U}_{\textsf{unique}}|$ is at most $\delta/2$. Since $(1-\frac{\epsilon}{2})\geq\frac{1}{1+\epsilon}$, so we have ${\mathsf{Pr}}\!\left[\mathsf{Fail};{{\mathsf{U}}_{1,\ldots,k}}\right]\leq\frac{\delta}{2}$.